OpenStudy (anonymous):

Please Help! Three times the measure of an angle is equal to 15 less than twice the measure of it's complement. the measure of each of the angles is 55 and 35. true or false?

OpenStudy (anonymous):

If you let "x" be the measure of the first angle, and realize that complementary angles add to "90", the complement is "90 - x". 3x = 2(90 - x) - 15 And when you solve for "x", you'll have your answer.

OpenStudy (anonymous):

What do you get for "x" ? @Frangg

OpenStudy (anonymous):

x = 33 False.

OpenStudy (anonymous):

This might help: a(b + c) = ab + ac

OpenStudy (anonymous):

3x = 2(90 - x) - 15 -> 3x = 180 - 2x - 15 3x = -2x + (180 - 15) -> 3x = -2x + 165 Now, add 2x to both sides: 5x = 165 -> x = 33 and its complement is 57

OpenStudy (anonymous):

All good now, @Frangg ?

OpenStudy (anonymous):

lmao yuuuuuuuup! thanks :D

OpenStudy (anonymous):

uw!

OpenStudy (anonymous):

how do we become fans of people?

OpenStudy (anonymous):

Click on the name to the right of the name's icon. Wait a moment and then click on the "Become a fan" button. Refer to the attachment.

OpenStudy (anonymous):

yayy i got it thanks! :)

OpenStudy (anonymous):

Your welcome.